An investor invested a total of $3,200 in two mutual funds. One fund earned a 4% profit while the other earned a 2% profit. If the investor's total profit was $98, how much was invested in a mutual fund?
Answer by mananth(16949)

part I 4.00% per annum ---x
part II 2.00% per annum ---y

x+y=3200 ------------------------1
4.00% x + 2.00% y= 98
Multiply by 100
4x+2y= 9800.00 --------2

Multiply (1) by -4
we get

-4x-4y= -12800.00

Add this to (2)

0x-2y= -3000

divide by -2

y = 1500 investment at 2.00%
Balance 1700 investment at 4.00%
CHECK
1500 --------- 2.00% ------- 30.00
1700 ------- 4.00% ------- 68.00
Total -------------- 98.00